# Exploiting Boosting in Hyperdimensional Computing for Enhanced Reliability in Healthcare

### v1

@@ I Introduction
Hyperdimensional computing (HDC) is a rapidly advancing field within artificial intelligence, offering numerous valuable qualities that apply to various areas. HDC’s ability to efficiently encode and process data in high-dimensional spaces has generated substantial interest, particularly in machine learning and data analysis [1, 2, 3]. As we contemplate the practical use of HDC-based models in real-world applications, we encounter crucial requirements beyond mere accuracy. These requirements encompass robustness, reliability, consistency, and resource efficiency [4]. Despite these requirements, analyses of HDC in high-dimensional spaces has not yet been fully explored which can lead to overfitting by selecting very high dimension above only performance [5]. Our research shows that the partitioning of high-dimensional spaces directly correlates with the utility of the space, fulfilling those requirements under certain condition.
@@ II-A Boosting methods on Machine Learning
To address the limitations of individual predictive models, researchers have explored various ensemble methods that seek to enhance performance by combining multiple models. One such successful ensemble technique is boosting, which sequentially integrates multiple simple models called weak learners and aggregates their predictions to make final decisions. These weak learners are trained to compensate for each other’s prediction errors and can take various forms, including shallow decision trees or a few layers of multi-layer perceptions. One of the pioneering works in boosting is Adaptive Boosting (AdaBoost) [8]. AdaBoost begins by assigning equal weight values to data points within a training set. It then sequentially trains each weak learner, focusing on identifying misclassified data points and increasing their weights to ensure subsequent weak learners prioritize their correct classification. Each weak learner is assigned a confidence value based on its prediction performance, and the final prediction is determined through a confidence-weighted voting mechanism among the weak learners. Another notable approach is the Gradient Boosting Machine (GBM) [11], which differs from AdaBoost by seeking to minimize the residual error rather than assigning weights to data points. GBM also follows a sequential training process, with each weak learner striving to reduce the residual error left by the previous model through gradient descent. Subsequent boosting algorithms, such as Light GBM [12] and XGBoost [13], adopted this residual error minimization approach due to its demonstrated effectiveness in comparison to the data-point weighting strategy initially employed by AdaBoost.
@@ II-B Healthcare with wearable device
The WESAD dataset[14] stands at the forefront of wearable stress and affect detection, providing a comprehensive collection of physiological and motion data from both wrist- and chest-worn devices in a controlled lab environment with 15 subjects. Distinguishing itself through diverse sensor modalities, WESAD captures data such as blood volume pulse, electrocardiogram, electrodermal activity, electromyogram, respiration, body temperature, and three-axis acceleration. Notably, the dataset introduces a unique dimension by including three affective states—neutral, stress, and amusement, enriching its applicability in studying stress and emotions. The prevalent state-of-the-art solution involves varied convolutional neural network (CNN) architectures[15]. Amidst this landscape, Hyperdimensional Computing (HDC) emerges as a compelling alternative to CNN models. The inherent resource constraints of wearable devices, exemplified by WESAD, align seamlessly with HDC’s capabilities. Offering lightweight and online learning, HDC proves to be a natural fit for the intricacies of the WESAD dataset. Unlike conventional CNNs, HDC excels in managing multimodal data, addressing the varied sensor modalities present in WESAD. The fusion of WESAD’s rich sensor modalities and affective states with the computational robustness of HDC holds great promise for real-world applications in wearable stress and affect detection, presenting WESAD as an ideal testing ground for harnessing the power of hyperdimensional computing.
@@ II-C Hyperdimensional Classification
$\delta(V_{1},V_{2})=\frac{V_{1}^{\dagger}V_{2}}{||V_{1}||||V_{2}||}$ (1)
HDC is a computational technique inspired by the neural processes in the brain, where data points are encoded into a high-dimensional space. The learning process in HDC involves extracting universal patterns that define each label in a dataset through matrix multiplication using Gaussian distribution values and trigonometric activation functions(Binding, Permutation), such as cosine and sine. The encoded hypervectors $\mathbf{H}$ are linearly combined to generate representations for each class (Bundling), known as class hypervectors ${\mathbf{C}_{l}}$. These ${\mathbf{C}_{l}}$ are represented as single points in the $R^{D}$ space, and upon end of training, a point is produced for each label in the classification task, as illustrated in Figure 5. During the inference phase of HDC, a query hypervector $\mathbf{H}$ is encoded, and its similarity to the ${\mathbf{C}_{l}}$ is calculated. The class whose $\mathbf{H}$ exhibits the highest similarity to the query is predicted as the output. Recently, the similarity computation between each data point and the ${\mathbf{C}_{l}}$ using a similarity function $\delta(\mathbf{H},\mathbf{C}_{l})$ has been employed to guide model updates  [4]. This is based on assessments of similarity and correctness, facilitating an iterative refinement process. Such a process enhances the HDC model’s ability to capture underlying data patterns, thereby improving its predictive accuracy.
  • (1) Bundling consists of an addition of multiple hypervectors into a single hypervector, $\vec{R}=\vec{V}_{1}+\vec{V}_{2}$. This process transforms into memorization in high-dimensional space.
  • (2) Binding associates multiple orthogonal hypervectors (e.g., $\vec{V}_{1}$, $\vec{V}_{2}$) into a single hypervector ($\vec{R}=\vec{V}_{1}*\vec{V}_{2}$). The binded hypervector is a new object in HDC space which is orthogonal to all input hypervectors ($\delta(\vec{R},\vec{V}_{1})\simeq 0$ and $\delta(\vec{R},\vec{V}_{2})\simeq 0$).
  • (3) Permutation defines as a single rotational shift. The permuted hypervector will be nearly orthogonal to its original hypervector ($\delta(\vec{V}_{1},\rho\vec{V}_{1})\simeq 0$).
@@ IV Evaluation
In our experimental setup, we employed a single GPU, the GeForce RTX 4070, and a 12th generation Intel(R) Core(TM) i7-12700K CPU. We conducted a total of 10 runs for each experiment, utilizing datasets that included WESAD[14], the Nurse Stress Dataset[17], and the Stress-Predict Dataset[18] to evaluate accuracy and time efficiency. For all other aspects of the experiments, we exclusively employed the WESAD[14] dataset. The WESAD[14] dataset holds a prominent position as a benchmark dataset tailored for stress detection research. It encompasses multimodal sensor data collected from wearable devices such as Empatica E4 and RespiBAN, sourced from 16 subjects. This dataset encompasses a rich array of physiological and motion data indicators, encompassing EDA, ECG, EMG, and BVP readings. The Stress-Predict Dataset[18] and Nurse Stress Dataset[17] datasets, on the other hand, are solely collected via the E4 sensor and share features like EDA, ECG, EMG, and BVP. Notably, the Nurse Stress Dataset[17] and Stress-Predict Dataset[18] datasets include 37 and 15 subjects, respectively. We performed stress level classification, reducing it to three labels (good, common, stress), and the test data was organized based on subject units, with all results grounded in the test dataset.
Datasets undergo preprocessing using a moving average filter with a window size of 30, extracting statistical measures like minimum, maximum, mean, and standard deviation. Due to the varied ranges of these statistics, normalization is applied to ensure consistent scaling. Our experimental models include AdaBoost with learning rate (lr) set to 1.0 and 10 estimators, Random Forest with bootstrap enabled and 10 estimators, XGBoost with 10 estimators, SVM with a linear kernel, DNN consists of convolutional layers with a learning rate of 0.001, 4 linear layers [2048,1024,512,classes], ReLU activation, and dropout. Additionally, OnlineHD is configured with dimension adjustment, bootstrap enabled, lr of 0.035, and a gaussian probability distribution (0, 1). These setups ensure a comprehensive evaluation of model performance across various algorithms in the experimental framework. In the ensemble models, we employed $N_{L}=10$. The HDC model was configured with $D_{total}$ values spanning the range ${10,20,\dots,1000,4000,10000}$. As for the BoostHD model, the weak learner ($wl$) dimensionality ($D_{wl}$) was set as $D_{total}/L_{N}$.
@@ IV-A Performance
In our comprehensive experimentation, we subjected the BoostHD algorithm to rigorous evaluation across three distinct datasets: WESAD[14], Nurse Stress Dataset[17], and Stress-Predict Dataset[18]. The focal points of our assessment encompassed key metrics such as accuracy, inference time, and training time. Our findings unveiled the remarkable prowess of BoostHD, consistently attaining the highest accuracy levels across all three datasets, as presented in Table I. Notably, when compared to OnlineHD on the WESAD[14] dataset, the accuracy achieved by BoostHD remained outside the range of two standard deviations. This distinction was further underscored by OnlineHD’s inability to reach a 98% accuracy threshold in any of the ten independent trials, while BoostHD consistently surpassed this benchmark, highlighting its superior predictive accuracy.
BoostHD strategically harnesses in-memory learning, epitomized by HDC, and augments it with adaptive learning mechanisms within an ensemble framework. This unique fusion results in a significant acceleration of the learning process, even when training is serialized. Empirical assessments conducted within a GPU environment unequivocally established BoostHD’s superiority over DNNs, demonstrating substantially enhanced processing speed across the three benchmark datasets. Moreover, in a CPU environment, BoostHD showcased remarkable computational efficiency, achieving processing speeds 26 times faster than its DNN counterparts. Furthermore, our optimization efforts fine-tuned the BoostHD algorithm for parallel processing, yielding substantial gains in inference efficiency (as detailed in Table II). This optimization proved particularly beneficial when handling the relatively extensive input vectors in the Nurse Stress Dataset[17] and Stress-Predict Dataset[18]. The outcome was a significant reduction in the time required for inferring data from the test dataset, firmly establishing BoostHD as the model of choice for maximizing time efficiency across all models evaluated, especially in the context of the Nurse Stress Dataset[17] and Stress-Predict Dataset[18].

### v2

@@ I Introduction
Hyperdimensional computing (HDC) is a rapidly advancing field within artificial intelligence, offering numerous valuable qualities that apply to various areas  [1, 2, 3, 4, 5, 6, 7]. HDC’s ability to efficiently encode and process data in high-dimensional spaces has generated substantial interest, particularly in machine learning and data analysis [8, 9, 10, 11, 12, 13, 14, 15, 16, 17]. As we contemplate the practical use of HDC-based models in real-world applications, we encounter crucial requirements beyond mere accuracy. These requirements encompass robustness, reliability, consistency, and resource efficiency [18]. Despite these requirements, analyses of HDC in high-dimensional spaces has not yet been fully explored which can lead to overfitting by selecting very high dimension above only performance [19]. Our research shows that the partitioning of high-dimensional spaces directly correlates with the utility of the space, fulfilling those requirements under certain condition.
@@ II-A Boosting methods on Machine Learning
To overcome the limitations of individual predictive models, ensemble methods such as boosting have been developed to improve performance by combining multiple weak learners, such as shallow decision trees or simple neural networks. Boosting aggregates the predictions of these learners, trained sequentially to correct each other’s errors. A notable example is Adaptive Boosting (AdaBoost)[22], which assigns equal initial weights to data points, adjusts weights to focus on misclassified points, and uses confidence-weighted voting for final predictions. Gradient Boosting Machine (GBM)[25], in contrast, minimizes residual errors through gradient descent instead of reweighting data points. Extensions like LightGBM [26] and XGBoost [27] adopt this residual minimization strategy, offering improved performance over AdaBoost.
@@ II-B Healthcare with wearable device
The WESAD dataset [28] is a benchmark for wearable stress and affect detection, comprising physiological and motion data from wrist- and chest-worn devices collected from 15 subjects in a controlled lab setting. It includes diverse sensor modalities such as blood volume pulse, electrocardiogram, electrodermal activity, respiration, temperature, and three-axis acceleration, alongside three affective states—neutral, stress, and amusement. While convolutional neural networks (CNNs) dominate as the state-of-the-art solution [29], Hyperdimensional Computing (HDC) offers a resource-efficient alternative better suited for the constraints of wearable devices. HDC’s lightweight, online learning approach effectively handles WESAD’s multimodal data, making it a promising method for real-world applications in stress and affect detection. The combination of WESAD’s rich modalities and HDC’s computational efficiency underscores the dataset’s potential as a testing ground for advanced wearable technologies.
@@ II-C Hyperdimensional Classification
HDC is inspired by neural processes in the brain, encoding data into a high-dimensional space. It learns universal patterns for each label by encoding data points as hypervectors $\mathbf{H}$ through matrix multiplication with Gaussian distribution values and trigonometric activation functions (Binding), such as sine and cosine. These hypervectors are combined (Bundling) to create class hypervectors $\mathbf{C}_{l}$, which represent each label in $R^{D}$ space. During inference, a query hypervector $\mathbf{H}$ is encoded and compared to class hypervectors using a similarity function $\delta(\mathbf{H},\mathbf{C}_{l})$. This process is widely applicable, as demonstrated in  [30, 31, 32, 33, 34, 35, 36, 37].
Model updates now leverage similarity assessments to iteratively refine predictions [18], enhancing the model’s ability to capture patterns and improve accuracy:
$\delta(V_{1},V_{2})=\frac{V_{1}^{\dagger}V_{2}}{\|V_{1}\|\|V_{2}\|}$ (1)
Key operations in HDC include:
  • Bundling: Combines hypervectors into a single vector ($\vec{R}=\vec{V}_{1}+\vec{V}_{2}$), enabling memorization in high-dimensional space.
  • Binding: Associates orthogonal hypervectors ($\vec{R}=\vec{V}_{1}*\vec{V}_{2}$), creating a new hypervector orthogonal to inputs ($\delta(\vec{R},\vec{V}_{1})\simeq 0$).
@@ IV Evaluation
In our experimental setup, we used a GeForce RTX 4070 GPU and a 12th generation Intel(R) Core(TM) i7-12700K CPU. Each experiment was conducted over 10 runs, employing datasets including WESAD [28], the Nurse Stress Dataset [39], and the Stress-Predict Dataset [40] to evaluate accuracy and time efficiency. For other aspects of the experiments, only the WESAD dataset was used. The WESAD dataset, a benchmark for stress detection research, features multimodal data from wearable devices like Empatica E4 and RespiBAN collected from 16 subjects. It includes physiological and motion indicators such as EDA, ECG, EMG, and BVP. The Nurse Stress Dataset [39] and Stress-Predict Dataset [40], collected via the E4 sensor, also include EDA, ECG, EMG, and BVP data from 37 and 15 subjects, respectively. Stress level classification was reduced to three labels: good, common, and stress. Test data was organized by subject units, and all results were based on the test dataset.
Datasets were preprocessed using a moving average filter with a window size of 30, extracting statistical features such as minimum, maximum, mean, and standard deviation. To address varying ranges, normalization was applied to ensure consistent scaling. The experimental models included AdaBoost (learning rate = 1.0, 10 estimators), Random Forest (bootstrap enabled, 10 estimators), XGBoost (10 estimators), SVM (linear kernel), and a DNN comprising convolutional layers with a learning rate of 0.001, four linear layers [2048, 1024, 512, classes], ReLU activation, and dropout. Additionally, OnlineHD was configured with dimensional adjustment, bootstrap enabled, a learning rate of 0.035, and Gaussian distribution $N(0,1)$. For ensemble models, $N_{L}=10$. The HDC model was evaluated with $D_{total}$ values ranging from 10 to 10,000, while for the BoostHD model, the weak learner dimensionality ($D_{wl}$) was set to $D_{total}/N_{L}$.
@@ IV-A Performance
In our comprehensive experimentation, we subjected the BoostHD algorithm to rigorous evaluation across three distinct datasets: WESAD[28], Nurse Stress Dataset[39], and Stress-Predict Dataset[40]. The focal points of our assessment encompassed key metrics such as accuracy, inference time, and training time. Our findings unveiled the remarkable prowess of BoostHD, consistently attaining the highest accuracy levels across all three datasets, as presented in Table I. Notably, when compared to OnlineHD on the WESAD[28] dataset, the accuracy achieved by BoostHD remained outside the range of two standard deviations. This distinction was further underscored by OnlineHD’s inability to reach a 98% accuracy threshold in any of the ten independent trials, while BoostHD consistently surpassed this benchmark.
BoostHD strategically combines in-memory learning, exemplified by HDC, with adaptive learning mechanisms within an ensemble framework, resulting in a significantly accelerated learning process, even when training is serialized. Empirical evaluations in a GPU environment demonstrated BoostHD’s clear superiority over DNNs, achieving substantially faster processing speeds across the three benchmark datasets. In a CPU environment, BoostHD achieved remarkable computational efficiency, processing data 26 times faster than DNN counterparts. Additionally, optimization efforts tailored BoostHD for parallel processing, greatly improving inference efficiency Table II. These enhancements were particularly effective when processing the relatively large input vectors in the Nurse Stress Dataset [39] and Stress-Predict Dataset [40], resulting in a significant reduction in inference time. This firmly establishes BoostHD as the most time-efficient model across all evaluated datasets, particularly for the Nurse Stress Dataset [39] and Stress-Predict Dataset [40].